OpenStudy (anonymous):

Divide the following polynomial (x^3 + y^3) / (x - y)

OpenStudy (anonymous):

D ( x3 + y3 ) = D ( 4 ) , D ( x3 ) + D ( y3 ) = D ( 4 ) , (Remember to use the chain rule on D ( y3 ) .) 3x2 + 3y2 y' = 0 , so that (Now solve for y' .) 3y2 y' = - 3x2 , and $ y' = \displaystyle{ - 3x^2 \over 3y^2 } = \displaystyle{ - x^2 \over y^2 } $ .